OpenStudy (anonymous):

how do I calculate how many moles are in 10 ml of hydrogen gas?

OpenStudy (somy):

at RTP or STP?

OpenStudy (anonymous):

im not sure im doing a lab and it says calculate the moles of hydrogen gas collected and I collected 10 ml of it

OpenStudy (anonymous):

Probably RTP if you did it in lab.

OpenStudy (anonymous):

I remember the equation is PV=nRT, but that's about it.

OpenStudy (anonymous):

you are solving for n, the number of moles. Plug in your pressure, etc. and solve for n. Should look like \[\frac{ PV }{ RT } =n\]

OpenStudy (somy):

yeah most probably its RTP so your RTP volume is \(\sf24~dm^3\) and the formula you can use to find mole of H is \(\sf your~volume = mole \times RTP~volume\) your volume of H = 10mL mole = X RTP volume = \(\sf24~dm^3\) so we need to find mole, thats why we gotta make mole of this formula as subject of this formula \[\sf mole = \frac{ your ~volume }{ RTP~volume }\] so use the formula and get the answer